Calculate Grxn at 298 K under the conditions shown below for the following reaction.              

                    2 H2S(g) + 3 O2(g)  2 SO2(g) + 2 H2O(g) G°rxn = ?

G°f (kJ/mol)    -33.4                      -300.1         -228.6

P(H2S) = 1.4 atm,   P(O2) = 0.5atm   P(SO2) = 2.2atm and P(H2O) = 0.8 atm
Answer #1

44F -33.千 - 300.1 - 228. 6 0 =0.8 atm roduct (2x-300.) + (2x-22.(x -33-4)+ o X n 12.4 = 12-64